# Exploiting Priors from 3D Diffusion Models for RGB-Based One-Shot View Planning

### v1

@@ Abstract
Object reconstruction is relevant for many autonomous robotic tasks that require interaction with the environment. A key challenge in such scenarios is planning view configurations to collect informative measurements for reconstructing an initially unknown object. One-shot view planning enables efficient data collection by predicting view configurations and planning the globally shortest path connecting all views at once. However, geometric priors about the object are required to conduct one-shot view planning. In this work, we propose a novel one-shot view planning approach that utilizes the powerful 3D generation capabilities of diffusion models as priors. By incorporating such geometric priors into our pipeline, we achieve effective one-shot view planning starting with only a single RGB image of the object to be reconstructed. Our planning experiments in simulation and real-world setups indicate that our approach balances well between object reconstruction quality and movement cost.
@@ I Introduction
Without any prior knowledge about the environment, a common strategy is to plan the next-best-view (NBV) iteratively based on the current reconstruction state [7, 33, 29, 37, 18, 19, 23]. However, NBV planning only generates a local path to the subsequent view and cannot effectively distribute the mission time or movement budget, resulting in suboptimal view planning performance. An alternative line of work considers one-shot view planning [26, 28, 6]. Given initial measurements of an object to be reconstructed, one-shot view planning predicts a set of views at once and computes the globally shortest path connecting them. A robot’s sensor then follows the planned path to collect measurements, which are used for object reconstruction after the data acquisition is completed. By decoupling data collection and object reconstruction, these approaches do not rely on iterative map updates for adaptive view planning online during a mission. To perform one-shot view planning, prior knowledge about the object is required. Previous works consider planning priors based on multi-view images or partial point cloud observations. However, they either only handle a fixed view configuration [28] or rely on depth sensors [6, 26].
@@
The main contribution of this work is a novel RGB-based one-shot view planning approach that exploits the geometric priors from 3D diffusion models. Our approach enables view planning with an object-specific view configuration for object reconstruction as shown in Fig. 1. A key component of our pipeline is a 3D diffusion model that outputs a 3D mesh of the object given one RGB image as input. This generated mesh is a proxy to the inaccessible ground truth 3D model and serves as the basis for our one-shot view planning. Given the generated 3D mesh, we convert the one-shot view planning into a customized set covering optimization to calculate the minimum set of views that densely covers the mesh, which we solve using linear programming. Our approach adaptively places the views and follows the globally shortest path for collecting informative RGB images around the object. After the data collection, we train a Neural Radiance Field (NeRF) using all collected images to acquire the object’s 3D representation.
To the best of our knowledge, our approach is the first to leverage 3D diffusion models for view planning. We make the following three claims: (i) we exploit the powerful 3D diffusion models to enable our one-shot view planning starting with only one RGB image as input; (ii) we cast the one-shot view planning as a customized set covering optimization problem, yielding view configurations suitable for RGB-based object reconstruction using NeRFs; (iii) our one-shot view planning allows for adaptive view placement to account for varying object geometries, achieving a better trade-off between movement cost and reconstruction quality compared to baselines. We conduct extensive experiments on publicly available object datasets and in real world scenarios, demonstrating the applicability and generalization ability of our approach. To support reproducibility and future research, our implementation will be open-sourced at: https://github.com/psc0628/DM-OSVP
@@ III-B One-Shot View Planning as Set Covering Optimization
One-shot view planning problem can be treated as a conventional set covering optimization. Since solving this optimization problem necessitates an explicit 3D representation of the object to be reconstructed, previous works [26, 6] rely on depth sensors to acquire initial 3D models of the object. Instead, by incorporating the geometric priors of 3D diffusion models into our planning pipeline, our approach solves the one-shot view planning problem in an RGB camera setup.
@@
Given and each , the conventional set covering optimization problem aims to find the minimum set of views required for completely covering the surface points. For instance, consider , , , and . The union of these three sets equals the entire surface set, i.e., . However, we can cover all surface points with only two sets, and . Vanilla set covering optimization requires that each surface point should be covered by at least one view. This definition aligns well with object reconstruction employing depth-sensing modalities [26, 27, 6], as surfaces can be recovered by direct depth fusion when provided with a corresponding point cloud observation. However, for RGB-based object reconstruction using NeRFs, map representation learning is achieved by minimizing the photometric loss when reprojecting hypothetical surface points back to 2D image planes, which requires that a surface point should be observed from different perspectives to recover its true 3D representation. This implies that planned views covering all surface points of the generated mesh once are not sufficient for object reconstruction using NeRFs.
@@ III-C Path Generation and Object Reconstruction
By planning all required views before data collection, the one-shot view planning paradigm shows a major advantage in reduced movement costs. Given the optimized set of views introduced above, we plan the globally shortest path connecting all views by solving the shortest Hamiltonian path problem on a graph, which is similar to the traveling salesman problem but without returning to the starting node [22]. The robot’s RGB camera follows the global path to acquire RGB measurements at planned views.
@@ IV-A Experimental Setup
In our simulation experiments, we consider an object-centric hemispherical view space with 144 uniformly distributed view candidates for view planning. We set the view space radius to 0.3 m. We test our approach on geometrically complex 3D object models from the HomebrewedDB dataset [10]. The test objects are shown in Fig. 5. We normalize all objects to fit into a bounding sphere with a radius of 0.1 m. All RGB measurements are at px px resolution. We adopt a grid size of in OctoMap for voxelizing the mesh surface points. The set covering optimization for view planning runs on an Intel i7-12700H CPU, while NeRFs training is conducted on an NVIDIA RTX3060 laptop GPU.
To evaluate NeRF reconstruction quality, we report the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) [20]. Additionally, we evaluate reconstruction efficiency in terms of inference time for view planning and accumulated movement cost for data collection in Euclidean distance.
@@ IV-B Analysis on Multi-View Constraints
In this section, we explore the influence of multi-view constraints introduced in Sec. III-B. We test our methods across varying values from to , as detailed in TABLE I. The outcomes reveal that: (1) with increasing values, our optimizer outputs on average more views for covering the mesh surfaces; (2) both PSNR and SSIM metrics exhibit a consistent improvement with increasing . Specifically, we achieve the highest performance gain by changing to , justifying our modification of the set covering optimization to account for RGB-based object reconstruction using NeRFs; (3) our method reaches its peak performance at the value of 6, while increasing to a higher value does not yield a statistically significant performance improvement.
@@ IV-D Evaluation of View Planning for Object Reconstruction
Comparison to Random Selection. As shown in Fig. 7, our RGB-based one-shot view planning approach surpasses the one-shot Random baseline across all values in terms of PSNR and SSIM. This is because heuristic random methods do not utilize any available information about the objects, in contrast to our approach. The Random method exhibits a slightly lower movement cost. We believe that this occurs since it can produce spatially clustered views, yielding poorer reconstruction quality. These findings confirm that leveraging powerful geometric priors from 3D diffusion models significantly benefits one-shot view planning for RGB-based object reconstruction.
@@
Comparison to PRV. Since the PRV method obtains the number of views by predicting the upper limits of NeRF representations, it is not associated with values and is represented by a dashed line in Fig. 7. The results indicate that the proposed RGB-based one-shot view planning approach, with an setting, delivers nearly identical quality metrics in PSNR and SSIM when compared to PRV, yet it benefits from reduced movement cost. Moreover, when is adjusted to 6, our method surpasses PRV in terms of PSNR and SSIM quality while still maintaining a slightly lower movement cost. This confirms that our adaptive object-specific view configuration is superior to fixed view configurations in PRV for handling varying geometries of objects.
In conclusion, our RGB-based one-shot view planning method demonstrates several advantages over the baselines. By integrating powerful geometric priors from 3D diffusion models, our method effectively leverages available object information, resulting in more informative and better distributed views. Moreover, our approach showcases superior adaptability through its adaptive view configuration mechanism. Unlike the fixed view configuration in PRV, our method dynamically adjusts the view configurations for different objects based on their varying geometries. However, we observe a longer inference time of our method compared to the PRV and random methods, primarily due to the constraints imposed by the generation process of the diffusion model (approximately s) and the online optimization process (approximately s). We plan to improve this in the future.
@@ IV-F Real-World Experiments
We deploy our approach in a real world tabletop environment using a UR5 robot arm with an Intel Realsense D435 camera mounted on its end-effector (only the RGB optical camera is activated). MoveIt [1] is employed for robotic motion planning. The accompanying video111https://youtu.be/EKZPHb5-UZk illustrates the online reconstruction process.
To validate our findings in Sec. IV-D, we compare our method against baselines in the real world. It is worth noting that due to imperfect camera poses and noise in real world experiments, the pose optimization functionality implemented in Instant-NGP is enabled during our NeRF training. The experimental environment and comparisons are shown in Fig. 9. From the results, we confirm that (1) our method generalizes to real world environments; and (2) our method adaptively plans view configurations according to object geometries to achieve higher PSNR with lower movement costs compared to the PRV and NBV methods. Our method achieves peak performance at , which is larger than the value of determined in Sec. IV-B. This might be caused by the noise in the camera pose and images, making it challenging for view planning tasks. We observe a similar slight performance reduction for the NBV methods.
@@ V Conclusions
In this paper, we present a novel one-shot view planning method starting with only a single RGB image of the unknown object to be reconstructed. The proposed method exploits priors from 3D diffusion models as a proxy to the inaccessible ground truth 3D model as the basis for one-shot view planning. We develop a customized variant of the set covering optimization problem tailored for NeRF-based reconstruction, which aims to adaptively compute a view configuration that densely covers the generated mesh from 3D diffusion models. We compute a globally shortest path on this view configuration, corresponding to the minimum travel distance. Our experiments validate that utilizing geometric priors from 3D diffusion models yields more informative views compared to random and next-best-view methods. When compared to the state-of-the-art RGB-based one-shot baseline, our adaptive view placement based on varying object geometries demonstrates better view planning performance compared to a fixed view configuration. The real world experiment suggests the applicability of our method.

### v2

@@ Abstract
Object reconstruction is relevant for many autonomous robotic tasks that require interaction with the environment. A key challenge in such scenarios is planning view configurations to collect informative measurements for reconstructing an initially unknown object. One-shot view planning enables efficient data collection by predicting view configurations and planning the globally shortest path connecting all views at once. However, prior knowledge about the object is required to conduct one-shot view planning. In this work, we propose a novel one-shot view planning approach that utilizes the powerful 3D generation capabilities of diffusion models as priors. By incorporating such geometric priors into our pipeline, we achieve effective one-shot view planning starting with only a single RGB image of the object to be reconstructed. Our planning experiments in simulation and real-world setups indicate that our approach balances well between object reconstruction quality and movement cost.
@@ I Introduction
Without any prior knowledge about the environment, a common strategy is to plan the next-best-view (NBV) iteratively based on the current reconstruction state [7, 33, 29, 37, 18, 19, 23]. However, NBV planning only generates a local path to the subsequent view and cannot effectively distribute the mission time or movement budget, resulting in suboptimal view planning performance. An alternative line of work considers one-shot view planning [26, 28, 6]. Given initial measurements of an object to be reconstructed, one-shot view planning predicts a set of views at once and computes the globally shortest path connecting them. A robot’s sensor then follows the planned path to collect measurements, which are used for object reconstruction after the data acquisition is completed. By decoupling data collection and object reconstruction, these approaches do not rely on iterative map updates for object-specific view planning online during a mission. To perform one-shot view planning, prior knowledge about the object is required. Previous works consider planning priors based on multi-view images or partial point cloud observations. However, they either only handle a fixed view configuration [28] or rely on depth sensors [6, 26].
@@
The main contribution of this work is a novel RGB-based one-shot view planning approach that exploits the geometric priors from 3D diffusion models. Our approach enables view planning with an object-specific view configuration for object reconstruction as shown in Fig. 1. A key component of our pipeline is a 3D diffusion model that outputs a 3D mesh of the object given one RGB image as input. This generated mesh is a proxy to the inaccessible ground truth 3D model and serves as the basis for our one-shot view planning. Given the generated 3D mesh, we convert the one-shot view planning into a customized set covering optimization problem to calculate the minimum set of views that densely covers the mesh, which we solve using linear programming. Our approach places the object-specific views and follows the globally shortest path for collecting informative RGB images around the object. After the data collection, we train a Neural Radiance Field (NeRF) using all collected images to acquire the object’s 3D representation.
To the best of our knowledge, our approach is the first to leverage 3D diffusion models for view planning. We make the following claims: (i) we exploit the powerful 3D diffusion models to enable our one-shot view planning starting with only one RGB image as input; (ii) we design the one-shot view planning as a customized set covering optimization problem, yielding view configurations suitable for RGB-based object reconstruction using NeRFs. We conduct extensive experiments on publicly available object datasets and in real world scenarios, demonstrating the applicability and generalization ability of our approach. Our one-shot view planning allows for object-specific view placement to account for varying object geometries, achieving a better trade-off between movement cost and reconstruction quality compared to baselines. To support reproducibility and future research, our implementation is open-sourced at: https://github.com/psc0628/DM-OSVP
@@ III-B One-Shot View Planning as Set Covering Optimization
One-shot view planning can be treated as a conventional set covering optimization problem. Since solving this optimization problem necessitates an explicit 3D representation of the object to be reconstructed, previous works [26, 6] rely on depth sensors to acquire initial 3D models of the object. Instead, by incorporating the geometric priors of 3D diffusion models into our planning pipeline, our approach solves the one-shot view planning problem in an RGB camera setup.
@@
Given and each , the conventional set covering optimization problem aims to find the minimum set of views required for completely covering the surface points. For instance, consider , , , and . The union of these three sets equals the entire surface set, i.e., . However, we can cover all surface points with only two sets, and . Vanilla set covering optimization problem requires that each surface point should be covered by at least one view. This definition aligns well with object reconstruction employing depth-sensing modalities [26, 27, 6], as surfaces can be recovered by direct depth fusion when provided with a corresponding point cloud observation. However, for RGB-based object reconstruction using NeRFs, map representation learning is achieved by minimizing the photometric loss when reprojecting hypothetical surface points back to 2D image planes, which requires that a surface point should be observed from different perspectives to recover its true 3D representation. This implies that planned views covering all surface points of the generated mesh once are not sufficient for object reconstruction using NeRFs.
@@ III-C Path Generation and Object Reconstruction
By planning all required views before data collection, the one-shot view planning paradigm shows a major advantage in reduced movement costs. Given the optimized set of views introduced above, we plan the globally shortest path connecting all views by solving the shortest Hamiltonian path problem on a graph, which is similar to the traveling salesman problem [22]. The robot’s RGB camera follows the global path to acquire RGB measurements at planned views. We follow the point-to-point local path planning method [27] to avoid collisions with the object.
@@ IV-A Experimental Setup
In our simulation experiments, we consider an object-centric hemispherical view space with 144 uniformly distributed view candidates for view planning [28]. We set the view space radius to 0.3 m. We test our approach on geometrically complex 3D object models from the HomebrewedDB dataset [10]. The test objects are shown in Fig. 5. We normalize all objects to fit into a bounding sphere with a radius of 0.1 m. All RGB measurements are at px px resolution. We adopt a grid size of in OctoMap for voxelizing the mesh surface points. The set covering optimization for view planning runs on an Intel i7-12700H CPU, while NeRFs training is conducted on an NVIDIA RTX3060 laptop GPU.
To evaluate NeRF reconstruction quality, we report the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) [20] on 100 uniformly distributed novel views [28]. Additionally, we evaluate reconstruction efficiency by inference time for view planning and accumulated movement cost for data collection in Euclidean distance.
@@ IV-B Analysis on Multi-View Constraints
In this section, we explore the influence of multi-view constraints introduced in Sec. III-B. We test our methods across varying values from to , as detailed in TABLE I. The outcomes reveal that: (1) with increasing values, our optimizer outputs on average more views for covering the mesh surfaces; (2) both PSNR and SSIM metrics exhibit a consistent improvement with increasing . Specifically, we achieve the highest performance gain by changing to , justifying our modification of the set covering optimization problem to account for RGB-based object reconstruction using NeRFs; (3) our method reaches its peak performance at the value of 6, while increasing to a higher value does not yield a statistically significant performance improvement.
@@ IV-D Evaluation of View Planning for Object Reconstruction
Comparison to Random Selection. As shown in Fig. 7, our RGB-based one-shot view planning approach surpasses the one-shot Random baseline across all values in terms of PSNR and SSIM. This is because the heuristic Random method does not utilize any available information about the objects, in contrast to our approach. The Random method exhibits a slightly lower movement cost. We believe that this occurs since it can produce spatially clustered views, yielding poorer reconstruction quality. These findings confirm that leveraging powerful geometric priors from 3D diffusion models significantly benefits one-shot view planning for RGB-based object reconstruction.
@@
Comparison to PRV. Since the PRV method obtains the number of views by predicting the upper limits of NeRF representations, it is not associated with values and is represented by a dashed line in Fig. 7. The results indicate that the proposed RGB-based one-shot view planning approach, with an setting, delivers nearly identical quality metrics in PSNR and SSIM when compared to PRV, yet it benefits from reduced movement cost. Moreover, when is adjusted to 6, our method surpasses PRV in terms of PSNR and SSIM quality while still maintaining a slightly lower movement cost. This confirms that our object-specific view configuration is superior to fixed view configurations in PRV for handling varying geometries of objects.
In conclusion, our RGB-based one-shot view planning method demonstrates several advantages over the baselines. By integrating powerful geometric priors from 3D diffusion models, our method effectively leverages available object information, resulting in more informative and better distributed views. Moreover, our approach showcases superior adaptability through its object-specific view configuration mechanism. Unlike the fixed view configuration in PRV, our method dynamically adjusts the view configurations for different objects based on their varying geometries. However, we observe a longer inference time of our method compared to the PRV and Random methods, primarily due to the constraints imposed by the generation process of the diffusion model (about s) and the online optimization process (about s). We plan to improve this in the future.
@@ IV-F Real-World Experiments
We deploy our approach in a real world tabletop environment using a UR5 robot arm with an Intel Realsense D435 camera mounted on its end-effector (only the RGB optical camera is activated). MoveIt [1] is employed for robotic motion planning. The accompanying video111https://youtu.be/EKZPHb5-UZk illustrates the online reconstruction process where .
To validate our findings in Sec. IV-D, we compare our method against baselines in the real world. It is worth noting that due to imperfect camera poses and noise in real world experiments, the pose optimization functionality implemented in Instant-NGP is enabled during our NeRF training. The experimental environment and comparisons are shown in Fig. 9. From the results, we confirm that (1) our method generalizes to real world environments; and (2) our method plans object-specific view configurations according to object geometries to achieve higher PSNR with lower movement costs compared to the PRV and NBV methods. Our method achieves peak performance at , which is larger than the value of determined in Sec. IV-B. This might be caused by the noise in the camera pose and images, making it challenging for view planning tasks. We observe a similar slight performance reduction for the NBV methods.
@@ V Conclusions
In this paper, we present a novel one-shot view planning method starting with only a single RGB image of the unknown object to be reconstructed. The proposed method exploits priors from 3D diffusion models as a proxy to the inaccessible ground truth 3D model as the basis for one-shot view planning. We develop a customized variant of the set covering optimization problem tailored for NeRF-based reconstruction, which aims to compute an object-specific view configuration that densely covers the generated mesh from 3D diffusion models. We compute a globally shortest path on this view configuration, corresponding to the minimum travel distance. Our experiments validate that utilizing geometric priors from 3D diffusion models yields more informative views compared to Random and next-best-view methods. When compared to the state-of-the-art RGB-based one-shot baseline, our view planning based on varying object geometries demonstrates better performance compared to a fixed view configuration. The real world experiment suggests the applicability of our method.